The high school band has 12 more girls than boys. The band has a total of 106 students. How many girls are in the band?

Answer by nerdybill(7384)

The high school band has 12 more girls than boys. The band has a total of 106 students. How many girls are in the band?
.
Let b = number of boys
then from "The high school band has 12 more girls than boys." we get:
b+12 = number of girls
.
b + b+12 = 106
2b+12 = 106
2b = 94
b = 47 (boys)
.
girls:
b+12 = 47+12 = 59